# The impact of stellar bars on quenching star formation: insights from a spatially resolved analysis in the local Universe

### v1

@@ 5 Discussion
The leading outcome of our work is the presence of an anti-correlating feature in the spatially resolved MS relations associated with an almost annular region that is hosting the bar. This pattern suggests that the typical MS bending observed at the high end, both in the spatially resolved relation (e.g., Abdurro’uf & Akiyama, 2017) and in the integrated one (e.g., Popesso et al., 2019a), can be due also to the presence of a stellar bar and not only by the bulge, as usually suggested (e.g., Morselli et al., 2019). However, it is still unclear which are the actual drivers of such a bending of the MS. For this reason, it could be crucial to investigate this correlation across different morphological structures (e.g., Pessa et al., 2022). The fact that the bar-hosting region is typically less star-forming and probably on the way to be quenched (it shows lower values of SFR with respect to what would be predicted by the scaling relation characterising the disk) goes in the direction of the inside-out quenching scenario, which is supposed to play an important role in driving galaxies below the MS (e.g., Tacchella et al., 2015, 2018; Abdurro’uf & Akiyama, 2018). This observed feature could also suggest a co-evolution trend of the bar structure and the inner bulge: the bar favours star formation quenching and, thus, makes the bulge grow.
@@
The quenching track in the spatially resolved MS could be more or less prominent: it is characterised by different slopes depending on the host galaxy (Figs. 2–3). This variability can be interpreted as a consequence of the different evolutionary stages in which the stellar bar is observed. More recently formed structures are probably still funnelling cold gas towards the inner regions of their hosts, thus resulting in higher values of star formation rate and in a less obvious anti-correlation feature (e.g., NGC 1566). On the other hand, during the late evolutionary stages of stellar bars, the central regions of their host galaxies are expected to be nearly devoid of fuel for further star formation, as a consequence of the previous rapid consumption of gas. This could potentially result in a more pronounced quenching track (e.g., NGC 4725), thereby also favouring the bulge growth.
@@
As already mentioned in Sect. 2.4, the environment in which galaxies reside is believed to influence their properties and evolution. The galaxies in our sample are located in quite different environments: field, loose groups and clusters. Among the various external processes, ram pressure stripping or gravitational interactions may have influenced the outskirts of some galaxies, so we cannot completely exclude the effect of the environment. However, the presence of the anti-correlation track in the bar region seems to be common to all six objects, although with different prominence. Therefore, we conclude that the environment is not the main driver of the feature found in our results. At most, it could be responsible for scatter or second-order variations.
@@
In the majority of galaxies in our sample, the nuclear regions are less star-forming than the spiral arms, except for the case of NGC 3351, which appears to be still active in forming stars in the very centre. A more obvious dual behaviour of the inner regions was pointed out by Enia et al. (2020) for their unbarred galaxies. In that case, basically half of the sample exhibits higher values of in the nucleus, and the other half turns out to be more quenched. Even though we do not have sufficient statistics to infer the actual difference, the fact that our barred galaxies are usually more quiescent in the central regions may suggest a possible effect of the bar presence. Indeed, this morphological structure is thought to be able to funnel gas both inwards and outwards and this can ignite starburst episodes, thus producing a rapid consumption of fuel for star formation and the onset of a quenching phase (e.g., Fraser-McKelvie et al., 2020a; Géron et al., 2021).
Similar results pointing to a lower star formation in the bar region were also obtained in the work of Pessa et al. (2021), based on spatially resolved SFR measurements from optical emission line maps, confirming the idea of ‘star formation deserts’ proposed by James & Percival (2018).
@@
In an attempt to capture a possible common trend in the star formation activity of barred galaxies, we compute the MS relation in the – plane considering the accepted grid cells of all six galaxies (see Fig. 4). This is also done to provide a fair comparison with the final Main Sequence relation found in Enia et al. (2020) (Eq. 3). Simply looking at the distribution of points, it is evident that different galaxies populate distinct portions of the plane, resulting in a quite broad relation, where the typical anti-correlation feature is barely discernible. This suggests that there is significant variability from one galaxy to another, as it has actually already been highlighted in the presentation of our results (see Sect. 4). To find the overall best-fit MS, we subdivide the mass density interval into 16 bins. Within each bin, we compute the median log and log, while the error is taken as the standard deviation of the points inside the bin. Besides, we fit our binned data points (black diamonds markers with associated error bars in Fig. 4) with a log-linear relation using emcee (Foreman-Mackey et al., 2013) and we obtain:
@@
where and dex. The global scatter of this final fit is dex. The values for the slope and the zero point of the best-fit relation are consistent, within the uncertainties, with the ones found by Enia et al. (2020), reported in Eq. 3. However, our relation is more scattered, indeed the scatter of the MS of unbarred galaxies turned out to be dex. This implies that there is greater variability in the properties of barred galaxies compared to the ones of unbarred objects, although these findings are based on limited statistics. Moreover, in Fig. 4, it is evident that the majority of data points are distributed below the MS of unbarred galaxies (grey dashed line). Hence, barred galaxies are generally less star-forming than their unbarred counterparts. This is also confirmed when looking at the integrated properties of the galaxies in our sample, compared with those in the work of Enia et al. (2020). At the same range of total stellar mass, between , the star formation rates of unbarred galaxies are between 1 and 4 yr-1, while barred galaxies show values of SFR well below 1 yr-1 (with the only exception of NGC 5236, see Table 1).
@@
As anticipated in Sect. 2, two out of six galaxies (namely NGC 1566 and NGC 4579) are characterised by low AGN activity in their nuclear regions. Since the magphys code does not account for a possible AGN contribution to the SED, we check its output by visually inspecting the SED fitting results obtained for the central pixels, where fluxes might be dominated by AGN emission. In Fig. 5 we show the outcome for one of the central pixels and for one of the outermost pixels of NGC 1566, one of the two galaxies hosting a LLAGN, and of NGC 5236, which has no nuclear activity, as a comparison. According to our results, the SED fit in most of the central pixels of NGC 1566 does not exhibit any peculiar feature that can be attributed to the influence of an AGN (as clearly visible in the example shown in Fig. 5). The only exception is for the pixel sampling the very centre of the galaxy, where the AGN contribution manifests as a partial suppression of the PAHs features. Hence, we ultimately excluded the very central pixel of both NGC 1566 and NGC 4579 from our analysis, to avoid any potential contamination. However, it is worth highlighting that this choice does not affect our results, since the nuclear region still remains properly sampled. Furthermore, looking at the SED fit in the pixels that are just slightly offset with respect to the very centre, we notice that there is no obvious residual in the – m range (covered by WISE and the IRAC camera on board of Spitzer), where the dusty torus emission of the AGN would have appeared (e.g., Fritz et al., 2006). The absence of a significant excess in this wavelength range with respect to the modelled continuum flux indicates that the resulting amount of star formation is sufficient to explain the observed SED. This is also confirmed by the fact that, comparing the SED fit in one of the central pixels of NGC 1566 with that of NGC 5236, there are no significant differences in the shape of the best-fit curve (left panels of Fig. 5). What has been said so far is valid for both NGC 1566 (upper panels of Fig. 5) and NGC 4579 and supports our initial choice to include LLAGN-hosting galaxies in our sample, assuming that the presence of a LLAGN would not significantly affect our SED modelling. Concerning NGC 4579, we have to underline the fact that, even though the LLAGN seems not to have a relevant impact on the SED fitting of the central pixels, the suppressed SFR detected in the region hosting the bar may also be due to a radio jet cocoon and radio-jet-driven outflows into the multiphase ISM, as found by Ogle et al. (2023). Additionally, comparing the SED fit in the innermost sub-kpc region of our galaxies with that obtained in one of the outer pixels, we notice that they have a similar shape, supporting the idea of a negligible contribution of the LLAGN to the galaxy emission.
@@
Nevertheless, many controversial results emerged from previous studies when exploring the connection between the bar structure and the presence of a potential AGN (see Combes, 2003, 2006; Jogee, 2006, for reviews). More recently, while Cisternas et al. (2013), Goulding et al. (2017), and Silva-Lima et al. (2022) claimed that bars are not among the main mechanisms driving AGN feeding, Cisternas et al. (2015) and Alonso et al. (2018) revealed that there could be an interplay between bars and AGN, supporting the idea of bar-induced nuclear activity. On the other way around, Zee et al. (2023) suggested that the evolution of AGN can regulate bars’ shape: these elongated structures can be destroyed by the growth of central mass concentration and black holes. Shedding light on these disparate results concerning the the joint evolution of bars, AGN and SF is still a highly debated work in progress.
@@ 6 Summary and conclusions
In this work, we study the spatially resolved Main Sequence relation of six nearly face-on barred spiral galaxies in the local Universe, with the aim of unveiling the relevance of possible bar-driven quenching effects. We take advantage of the publicly available DustPedia archive to collect photometric information from the UV to the FIR for the entire sample. Then, a panchromatic SED fitting procedure is performed on sub-kpc scales, making use of the magphys code. We produce maps of stellar mass and SFR surface densities and relate them in the – plane.
We outline our key findings below:
All the selected galaxies show a peculiar feature in the resolved MS relation: the bar hosting region tends to be offset below the best-fit MS, whose slope is largely determined by the dominant disc structure. Indeed, a trace of anti-correlation in correspondence with the annular region that hosts the bar is clearly visible in the MS of every galaxy and this hints at a suppressed SFR, which could be associated with the onset of a quiescent phase. We argue that stellar bars are mostly responsible for depleting gas reservoirs, thus halting star formation in the central regions. This supports the inside-out quenching scenario and the idea that bars may affect the typically observed bending of the MS at high .
@@
The innermost central regions, typically associated with a galactic bulge structure, are characterised by values of SFR that are lower than what would be predicted by the best-fit MS, with the exception of NGC 3351.
Comparing our results with the ones obtained for a sample of normal spiral galaxies, we notice that barred objects are typically less star-forming than their unbarred counterparts, even though the slope and the zero point of our total MS are consistent within the uncertainties.
In future follow-up works, it would be useful to improve the statistics by selecting a larger sample of barred galaxies in order to further investigate the role of bar quenching and to seek the actual causal relationship between bar formation and the central ‘star formation desert’. In particular, it may be interesting to include barred and unbarred galaxies hosting AGN to investigate if any systematic difference is found compared to galaxies with no strong AGN activity. Moreover, high-resolution simulations (e.g., Illustris TNG) might be crucial in this kind of study to test the reproducibility of the observed quenching trend and to reconstruct the history of bar formation and the evolution of SFR. In this way, one should also be able to better constrain the actual impact of stellar bars on the suppression of SFR by quantifying this quenching mechanism over time from a cosmological perspective.
@@ Appendix A SFR = SFRUV + SFRIR vs SFRMAGPHYS
In this section, we present the comparison between the SFRs derived as the sum of the two empirical relations (Eqs. 1–2) and the ones obtained from magphys. Similarly to what has been done in the work of Enia et al. (2020), we tested the consistency of these values of SFR producing the plot in Figure 6. The scatter of the relation, represented by the red-scale contours, is given by the median absolute deviation and it has a value of 0.15 dex, which is small. The fact that this tight distribution of points is compatible with the 1:1 relation (black line) ensures that the use of SFRs computed as the sum of Eqs. 1 and 2, instead of those given by magphys, do not change the outcomes of this work.

### v2

@@ 5 Discussion
The leading outcome of our work is the discovery of an anti-correlating feature in the spatially resolved MS relations associated with an almost annular region that is hosting the bar. This pattern suggests that the typical MS bending observed at the high end, both in the spatially resolved relation (e.g. Abdurro’uf & Akiyama, 2017) and in the integrated one (e.g. Popesso et al., 2019a), could be due to the presence of a stellar bar and not only the bulge, as usually suggested (e.g. Morselli et al., 2019). However, the precise drivers of this bending of the MS remain unknown. For this reason, it could be crucial to investigate this correlation across different morphological structures (e.g. Pessa et al., 2022). The fact that the bar-hosting region is typically more weakly star forming and is probably on the way to being quenched (it shows lower values of SFR with respect to those predicted by the scaling relation characterising the disc) supports the inside-out quenching scenario, which is supposed to play an important role in driving galaxies below the MS (e.g. Tacchella et al., 2015, 2018; Abdurro’uf & Akiyama, 2018). This observed feature could also suggest co-evolution of the bar structure and the inner bulge: the bar favours star formation quenching and thus induces bulge growth.
@@
The quenching track in the spatially resolved MS can be more or less prominent: it is characterised by different slopes depending on the host galaxy (Figs. 2–3). This variability can be interpreted as a consequence of the different evolutionary stages in which the stellar bar is observed. More recently formed structures are probably still funnelling cold gas towards the inner regions of their hosts, resulting in higher SFRs and a less obvious anti-correlation feature (e.g. NGC 1566). On the other hand, during the late evolutionary stages of stellar bars, the central regions of their host galaxies are expected to be nearly devoid of fuel for further star formation as a consequence of the previous rapid consumption of gas. This could potentially result in a more pronounced quenching track (e.g. NGC 4725), thereby also favouring bulge growth.
@@
As mentioned in Sect. 2.4, the environment in which galaxies reside is believed to influence their properties and evolution. The galaxies in our sample are located in quite different environments: field, loose groups, and clusters. Among the various external processes, ram pressure stripping or gravitational interactions may have influenced the outskirts of some galaxies, and so we cannot completely exclude environmental effects. However, the presence of the anti-correlation track in the bar region seems to be common to all six objects, albeit with different levels of prominence. Therefore, we conclude that the environment is not the main driver of the feature found in our results; at most, it could be responsible for scatter or second-order variations.
@@
In the majority of the galaxies of our sample, the nuclear regions are less star forming than the spiral arms, except for the case of NGC 3351, which appears to be still active in forming stars in the very centre. A more obvious dual behaviour of the inner regions was pointed out by Enia et al. (2020) for their unbarred galaxies; approximately half of their sample exhibits higher values of in the nucleus, and the other half are more quenched. Although we do not have sufficient statistics to infer the actual difference, the fact that our barred galaxies are usually more quiescent in the central regions suggests a possible effect of the presence of the bar. Indeed, this morphological structure is thought to be able to funnel gas both inwards and outwards and this can ignite starburst episodes, thus producing a rapid consumption of fuel for star formation and the onset of a quenching phase (e.g. Fraser-McKelvie et al., 2020a; Géron et al., 2021).
@@
Similar results pointing to lower levels of star formation in the bar region were also obtained by Pessa et al. (2021) based on spatially resolved SFR measurements from optical emission line maps, confirming the idea of ‘star formation deserts’ proposed by James & Percival (2018). In an attempt to capture a possible common trend in the star formation activity of barred galaxies, we computed the MS relation in the – plane, considering the accepted grid cells of all six galaxies (see Fig. 4). This is also done to provide a fair comparison with the final MS relation found in Enia et al. (2020) (Eq. 3). Simply looking at the distribution of points, it is evident that different galaxies populate distinct portions of the plane, resulting in a quite broad relation, where the typical anti-correlation feature is barely discernible. This suggests that there is significant variability from one galaxy to another, as is highlighted in the presentation of our results (see Sect. 4). To find the overall best-fit MS, we subdivided the mass-density interval into 16 bins. Within each bin, we computed the median log and log, while the error is taken as the standard deviation of the points inside the bin. In addition, we fitted our binned data points (black diamonds markers with associated error bars in Fig. 4) with a log-linear relation using emcee (Foreman-Mackey et al., 2013) and obtain:
@@
where and dex. The global scatter of this final fit is dex. The values for the slope and the zero point of the best-fit relation are consistent —within the uncertainties— with those found by Enia et al. (2020), reported in Eq. 3. However, our relation is more scattered; indeed they found the scatter of the MS of unbarred galaxies to be dex. This implies that there is greater variability in the properties of barred galaxies compared to those of unbarred objects, although these findings are based on limited statistics. Moreover, in Fig. 4, it is evident that the majority of data points are distributed below the MS of unbarred galaxies (grey dashed line). Hence, barred galaxies are generally less star forming than their unbarred counterparts. This is also confirmed when looking at the integrated properties of the galaxies in our sample and comparing them with those in the work of Enia et al. (2020). Within the same range of total stellar mass, that is, within , the SFRs of unbarred galaxies are between 1 and 4 yr-1, while barred galaxies show values of SFR well below 1 yr-1 (with the only exception being NGC 5236; see Table 1).
@@
As anticipated in Sect. 2, two out of six galaxies (namely NGC 1566 and NGC 4579) are characterised by low AGN activity in their nuclear regions. As the magphys code does not account for a possible AGN contribution to the SED, we checked its output by visually inspecting the SED fitting results obtained for the central pixels, where fluxes might be dominated by AGN emission. In Fig. 5 we show the outcome for one of the central pixels and for one of the outermost pixels of NGC 1566, one of the two galaxies hosting a LLAGN, and of NGC 5236, which has no nuclear activity, for comparison. According to our results, the SED fit in most of the central pixels of NGC 1566 does not exhibit any peculiar feature that can be attributed to the influence of an AGN (as clearly visible in the example shown in Fig. 5). The only exception is for the pixel sampling the very centre of the galaxy, where the AGN contribution manifests as a partial suppression of the PAHs features. Hence, we ultimately excluded the very central pixel of both NGC 1566 and NGC 4579 from our analysis, to avoid any potential contamination. However, it is worth highlighting that this choice does not affect our results, as the nuclear region still remains properly sampled. Furthermore, looking at the SED fit in the pixels that are just slightly offset with respect to the very centre, we notice that there is no obvious residual in the – m range (covered by WISE and the IRAC camera on board Spitzer), where the dusty torus emission of the AGN would have appeared (e.g. Fritz et al., 2006). The absence of significant excess in this wavelength range with respect to the modelled continuum flux indicates that the resulting star formation is sufficient to explain the observed SED. This is also confirmed by the fact that, comparing the SED fit in one of the central pixels of NGC 1566 with that of NGC 5236, there are no significant differences in the shape of the best-fit curve (left panels of Fig. 5). The conclusions listed above are valid for both NGC 1566 (upper panels of Fig. 5) and NGC 4579 and support our initial decision to include LLAGN-hosting galaxies in our sample, assuming that the presence of a LLAGN would not significantly affect our SED modelling. Concerning NGC 4579, we have to underline the fact that, although the LLAGN seems not to have a relevant impact on the SED fitting of the central pixels, the suppressed SFR detected in the region hosting the bar may also be due to a radio jet cocoon and radio-jet-driven outflows into the multi-phase ISM, as found by Ogle et al. (2023). Additionally, comparing the SED fit in the innermost subkpc region of our galaxies with that obtained in one of the outer pixels, we note that they have a similar shape, supporting the idea that the LLAGN makes a negligible contribution to the galaxy emission.
@@
Nevertheless, many controversial results emerged from previous studies when exploring the connection between the bar structure and the presence of a potential AGN (see Combes, 2003, 2006; Jogee, 2006, for reviews). More recently, while Cisternas et al. (2013), Goulding et al. (2017), and Silva-Lima et al. (2022) claimed that bars are not among the main mechanisms driving AGN feeding, Cisternas et al. (2015) and Alonso et al. (2018) revealed that there could be an interplay between bars and AGN, supporting the idea of bar-induced nuclear activity. Alternatively, Zee et al. (2023) suggested that AGN evolution can regulate the shape of a bar: these elongated structures can be destroyed by the growth of a central mass concentration or a black hole. Interpretation of these disparate results concerning the joint evolution of bars, AGN, and star formation is still a highly active area of research.
@@ 6 Summary and conclusions
In this work, we study the spatially resolved MS relation for six almost face-on barred spiral galaxies in the local Universe, with the aim of unveiling the relevance of possible bar-driven quenching effects. We took advantage of the publicly available DustPedia archive to collect photometric information on UV to FIR wavelengths for the entire sample. We performed a panchromatic SED fitting procedure on subkpc scales using the magphys code. We produced maps of stellar mass and SFR surface density and compared them in the – plane.
Our key findings can be summarised as follows:
All the selected galaxies show a peculiar feature in the resolved MS relation: the bar-hosting region tends to be offset below the best-fit MS, whose slope is largely determined by the dominant disc structure. Indeed, a slight anti-correlation in correspondence with the annular region that hosts the bar is clearly visible in the MS of every galaxy and this suggests a suppressed SFR, which could be associated with the onset of a quiescent phase. We argue that stellar bars are mostly responsible for depleting gas reservoirs, thus halting star formation in the central regions. This supports the inside-out quenching scenario and the idea that bars may affect the typically observed bending of the MS at high .
@@
The innermost central regions typically associated with a galactic bulge structure are characterised by lower SFR values than those predicted by the best-fit MS, with the exception of NGC 3351.
Comparing our results with those obtained for a sample of normal spiral galaxies, we note that barred objects are typically less actively star forming than their unbarred counterparts, even though the slope and the zero point of our total MS are consistent within the uncertainties.
In future follow-up works, it would be useful to improve the statistics by selecting a larger sample of barred galaxies in order to further investigate the role of bar quenching and to elucidate a possible causal relationship between bar formation and the central ‘star-formation desert’. In particular, it may be interesting to include barred and unbarred galaxies hosting AGN to investigate whether or not a systematic difference is found compared to galaxies with no strong AGN activity. Moreover, high-resolution simulations (e.g. Illustris TNG) might be crucial in this kind of study in order to test the reproducibility of the observed quenching trend and to reconstruct the history of bar formation and the evolution of SFR. In this way, one should also be able to better constrain the actual impact of stellar bars on the suppression of SFR by quantifying this quenching mechanism over time from a cosmological perspective.
@@ Acknowledgements.
- Sánchez-García et al. (2023) Sánchez-García, O., Cervantes Sodi, B., Fritz, J., et al. 2023, ApJ, 945, 99
@@ Appendix A SFR = SFRUV + SFRIR vs SFRMAGPHYS
In this section, we present a comparison between the SFRs derived as the sum of the two empirical relations (Eqs. 1–2) and the ones obtained from magphys. Similarly to what has been done in the work of Enia et al. (2020), we tested the consistency of the values of SFR producing the plot in Figure 6. The scatter of the relation, represented by the red-scale contours, is given by the median absolute deviation and has a value of 0.15 dex, which is small. The fact that this tight distribution of points is compatible with a 1:1 relation (black line) ensures that the use of SFRs computed as the sum of Eqs. 1 and 2, instead of those given by magphys, does not change the outcomes of this work.